I need help with dosage homework. Thank you!

At 0800 the client has a continuous heparin infusion running at 12 mL per hr. According to the hospital's heparin protocol, the heparin infusion should be increased by 100 units per hour each hour. The bag hanging contains heparin 40,000 units in D5W 250 mL. Calculate the new milliliter per hour flow rate for this heparin infusion. (Round to tenths).
Answer #1

The answer is 12.625 ml per Hour in order to increase 100 unit of Heparin.

Since 250 ml contains 40000 units of heparin, 1 ml contains 160 units . So 12 ml which is the current infusion volume which will contains 1920 units of Heparin.

As per the protocol 100 units to be increased per hour.

So the calculation would be as follows;

1920 +100 = 2020.

X = 2020/160 = 12.625 ml per hour.